A firm is employing 100 workers at a wage rate(W) = $l5lhour and 50 units of capital at a

rental rate® = $30lhour. At the firm's current input use, the marginal product of labor is 45 and

the marginal product of capital is 60. Is the firm producing its current level of output at the

minimum cost

a) Yes because MPL/W > MPK/R.

b) Yes because W<R

c) No because MPL/W > MPK/R

d) No because MPL < MPK


1
Expert's answer
2022-01-19T05:59:33-0500

"\frac{100}{15}\cdot 45>\frac{100}{30}\cdot 60,"

c) No because MPL/W > MPK/R.